OpenStudy (anonymous):

88x-5y=39 -8x + 3y =-1

OpenStudy (anonymous):

x= is -29/32 and -63/4

OpenStudy (anonymous):

how you got to this answer & thx

OpenStudy (anonymous):

wait is it two or one equation?

OpenStudy (anonymous):

oh it's one never mind let me recalculate

OpenStudy (anonymous):

88x-5y-39=0 y=88x/5-39/5 integer solution would be x=5n+3, y=88n+45 so, y=1/5(88x-3) =)

OpenStudy (anonymous):

thx alot